# HawkRover: An Autonomous mmWave Vehicular Communication Testbed with Multi-sensor Fusion and Deep Learning

### 03 Jan 2024 v1

@@ Abstract
Connected and automated vehicles (CAVs) have become a transformative technology that can change our daily life. Currently, millimeter-wave (mmWave) bands are identified as the promising CAV connectivity solution. While it can provide high data rate, their realization faces many challenges such as high attenuation during mmWave signal propagation and mobility management. Existing solution has to initiate pilot signal to measure channel information, then apply signal processing to calculate the best narrow beam towards the receiver end to guarantee sufficient signal power. This process takes significant overhead and time, hence not suitable for vehicles. In this study, we propose an autonomous and low-cost testbed to collect extensive co-located mmWave signal and other sensors data such as LiDAR (Light Detection and Ranging), cameras, ultrasonic, etc, traditionally for “automated”, to facilitate mmWave vehicular communications. Intuitively, these sensors can build a 3D map around the vehicle and signal propagation path can be estimated, eliminating iterative the process via pilot signals. This multimodal data fusion, together with AI, is expected to bring significant advances in “connected” research.
@@ I Introduction
Along this direction, to date, a few works have explored to using the out-of-band information for facilitating mmWave transmission. For example, in [5], the authors used cameras to localize receiver and then guide the transmitter to use direct beamforming towards the receiver. This approach, however, will not likely work well for non-line-of-sight (NLoS) scenarios. Then in [6], the proposed multi-sensor fusion that consists of input from radar and camera and have shown an improved performance even under NLoS. In [7], a field measurement dataset is constructed, which contains LiDAR, camera, radar, and mmWave signal directions. The tested scenario is also versatile, from vehicle (dynamic) to pedestrians (relative static). However, existing testbeds suffer from high cost and are difficult to configure and use, which limits their potential
@@
  • The proposed HawkRover consists of commercial-off-the-shelf (COTS) TP-Link Talon AD7200 routers and a RC car, which provides a cost-effective solution for obtaining mmWave wireless and co-located multi-sensory data. All data are time-synchronized through Internet time.
@@
  • Our results show that the constructed dataset and the DL algorithm perform good in real-world data, proving its potential in future V2X mmWave networks.
@@ II-A System overview and implementation
The HawkRover system architect is shown in Fig. 1. This overall design is versatile, allowing for adding sensors in a plug-and-play fashion. In particular, we use ROS (Robot Operating System) as the main software platform. ROS is a middleware between real operating system (Linux) and hardware. It contains a set of software libraries that are optimized to interact with sensor devices and has being widely adopted in robot community [8]. Docker is a virtualization suite and can dynamically allocate ROS system resources to each sensor, hence allows for better utilization on limited onboard resources. ROS will collect data from cameras, LiDAR, inertial measurement unit (IMU), and mmWave modules. When generating data, all sensors retrieve global synchronized time, which is critical for fusion algorithm. For indoor tests, data from cameras, LiDAR and IMU will also be used to enable autonomous driving on RC car, so this platform can operate without intervention. Lastly, since huge amount of data will be generated, external storage unit is deployed through wireless connection. Efficient data transmission between ROS and storage unit is enabled by message queuing telemetry transport (MQTT), a lightweight, loss-less, and bi-directional network layer protocol. Below is a brief list of main tools used in this HawkRover.
@@
ROS allows publisher/subscriber mode where each sensor interface acts as a publisher to push generated data to the subscriber (Linux), a timestamp is also recorded when data are admitted. Second, data will be temporarily stored at onboard device then enters MQTT pipeline, which connects to a much larger external storage unit. The mmWave device we choose is not only cost and power efficient, but also has a simple beamforming mechanism (pre-defined codebook and exhaustive beam sweep), which is likely to be adopted in future vehicular communication. A complete set of indoor package costs around \$1,300. This is possible with the low-cost mmWave router and RC car. Our setup is much lower than existing solutions. For example, one pair of mmWave TX and RX unit from universal software radio peripheral (USRP), each with flexible number of antennas, costs more than \$200,000 [9]. The cost of [7] can exceed \$10,000 due to higher end mmWave devices.
@@
As mentioned before, the proposed testbed can be operated autonomously, thanks to the “self-driving” capability enabled by Donkey car [10]. It utilizes the front camera to identify the edges of the road and continuously control the steering angles and throttles. In the lab, we have setup a large canvas with dark tapes acting as the “road”. Fig. 2 shows the prototype of HawkRover.
@@ III-B Beam prediction with multimodal sensor fusion
The aim is to directly predict best pair of codebooks from multisensor input, skipping channel estimation, as well as exhaustive search stage. Since enormous datasets can be generated from the HawkRover platform, our specific goal is to directly find the mapping:
$(b_{UE}^{*},b_{BS}^{*})=\max_{b_{UE},b_{BS}}f(\mathcal{C,L,I,P})$ (1)
@@
In Fig. 5, the overall structure of fusion algorithm is presented. At high-level, each sensor’s input is processed through convolutional neural network (CNN) and recurrent neural network (RNN). The former is usually used to process image, while the latter is for sequential data, perfectly fit for numerical wireless and location data. A fusion block is applied to concatenate each extracted feature, and then all features feed into LSTM block for temporal correlations. Lastly, the softmax-based output layer is utilized to determine the beam index with highest probability.
In CNN, convolution (Conv), rectified linear unit (ReLU), and pooling are three fundamental building blocks. Each serves a distinct function. Conv layers perform a mathematical operation that takes two inputs, such as an image matrix and a filter or kernel, and produces a feature map. This operation involves sliding the filter over the image and computing the dot product at each position. Convolution preserves the spatial relationship between pixels by learning image features using small squares of input data. The ReLU function applies a non-linear activation to the output of the convolution operation. It replaces all negative pixel values in the feature map with zero. This non-linearity introduces the ability to solve more complex problems by allowing the network to create complex mappings from the input to the output, improving the network’s ability to learn from the data. Pooling reduces the dimensionality of each feature map while retaining the most important information. Max pooling, for example, slides a window over the input and takes the maximum value at each step. By doing so, pooling layers reduce the number of parameters, lessen computational complexity, and provide a form of translation invariance. They help the network to be more robust to the exact location of features in the input space.
@@ IV A Case Study and Results
In this section, we provide a case study for one particular dataset collected in the summer of 2022. The location is a typical lab space. Note that during data collection, normal daily activities are not restricted, for example, people walking in the room, lab furniture movements, etc. This, in turn, shows the robustness and realism of the dataset.
@@
The collected data is divided into two parts: training (80%) and testing (20%). Both training process and test experiments are implemented in PyTorch. As shown in Fig. 7, the optimal beamforming prediction result based on sensor fusion algorithm is presented. In DL, Top-$K$ is often used as an evaluation metric to assess the performance. For example, in our case, top-1 accuracy means the model’s most probable prediction must be correct for it to count as a success. Meanwhile, top-5 accuracy means any of the model’s five most probable predictions can be correct. We showcase the results from top-1 to top-5, each with three different setting: LiDAR only, LiDAR and camera fusion, and LiDAR, Camera, and IMU position fusion.
@@
We need to note that the mmWave spatial fingerprints of the lab room may have also contributed to this performance. As shown in our early works [12, 14, 13], mmWave signal can have a spatial resolution to sub-10 cm, and the DL may have learnt this implicit relationship of location and optimal beam directions. Nevertheless, sensor fusion approach can significantly reduce the beam tracking overhead, which is a very desired feature for vehicular scenarios.
@@ V Conclusions
mmWave bands are emerging as a viable solution for CAV communication. However, their high attenuation and mobility management pose considerable challenges. Our study introduces HawkRover, an autonomous and cost-effective testbed for collecting comprehensive mmWave and sensor data such as from LiDAR, cameras, and IMUs, to enhance vehicular communication. By DL-based sensor fusion, our method can predict signal paths more efficiently than iterative pilot signal processes. A case study for indoor scenario is given, which promises substantial progress in CAV communication research. In the future, asynchronous state update will be developed to preserve more sensor data and improve the prediction accuracy.

### 04 Jan 2024 v2

@@ Abstract
Connected and automated vehicles (CAVs) have become a transformative technology that can change our daily life. Currently, millimeter-wave (mmWave) bands are identified as the promising CAV connectivity solution. While it can provide high data rate, their realization faces many challenges such as high attenuation during mmWave signal propagation and mobility management. Existing solution has to initiate pilot signal to measure channel information, then apply signal processing to calculate the best narrow beam towards the receiver end to guarantee sufficient signal power. This process takes significant overhead and time, hence not suitable for CAVs. In this study, we propose an autonomous and low-cost testbed to collect extensive co-located mmWave signal and other sensor data such as LiDAR (Light Detection and Ranging), cameras, ultrasonic, etc, traditionally for “automated”, to facilitate mmWave vehicular communications. Intuitively, these sensors can build a 3D map around the vehicle and signal propagation path can be estimated, eliminating the iterative process via pilot signals. This multimodal data fusion, together with AI, is expected to bring significant advances in “connected” research.
@@ I Introduction
Along this direction, to date, a few works have explored to using the out-of-band information for facilitating mmWave transmission. For example, in [5], the authors used cameras to localize receiver and then guide the transmitter to use direct beamforming towards the receiver. This approach, however, will not likely work well for non-line-of-sight (NLoS) scenarios. Then in [6], the proposed multi-sensor fusion that consists of input from radar and camera has shown an improved performance even under NLoS. In [7], a field measurement dataset is constructed, which contains LiDAR, camera, radar, and mmWave signal directions. The tested scenario is also versatile, from vehicle (dynamic) to pedestrians (relative static). However, existing testbeds suffer from high cost and are difficult to configure and use, which limits their potentials.
@@
  • The proposed HawkRover consists of commercial-off-the-shelf (COTS) TP-Link Talon AD7200 routers and a RC car. This provides a cost-effective solution for obtaining mmWave wireless and co-located multi-sensory data. All the data are time-synchronized through Internet time.
@@
  • Our results show that the constructed dataset and the DL algorithm perform well in real-world data, proving their potential in future V2X mmWave networks.
@@ II-A System overview and implementation
The HawkRover system architect is shown in Fig. 1. This overall design is versatile, allowing for adding sensors in a plug-and-play fashion. In particular, we use ROS (Robot Operating System) as the main software platform. ROS is a middleware between real operating system (Linux) and hardware. It contains a set of software libraries that are optimized to interact with sensor devices and has been widely adopted in robot community [8]. Docker is a virtualization suite and can dynamically allocate ROS system resources to each sensor, hence allowing for better utilization on limited onboard resources. ROS will collect data from cameras, LiDAR, inertial measurement unit (IMU), and mmWave modules. When generating data, all sensors retrieve global synchronized time, which is critical for fusion algorithm. For indoor tests, data from cameras, LiDAR and IMU will also be used to enable autonomous driving on RC car, so that this platform can operate without intervention. Lastly, since a huge amount of data will be generated, an external storage unit is deployed through wireless connection. Efficient data transmission between ROS and storage unit is enabled by message queuing telemetry transport (MQTT), a lightweight, loss-less, and bi-directional network layer protocol. Below is a brief list of main tools used in this HawkRover.
@@
ROS allows the publisher/subscriber mode, where each sensor interface acts as a publisher to push the generated data to the subscriber (Linux). A timestamp is also recorded when data are admitted. The data will be temporarily stored on an onboard device before entering the MQTT pipeline, which connects to a much larger external storage unit. The mmWave device we choose is not only cost effective and power efficient, but also has a simple beamforming mechanism (pre-defined codebook and exhaustive beam sweep), which is likely to be adopted in future vehicular communication. A complete set of indoor package costs around \$1,300. This is possible with the low-cost mmWave router and RC car. Our setup is much lower than existing solutions. For example, one pair of mmWave TX and RX unit from universal software radio peripheral (USRP), each equipped with a flexible number of antennas, can easily cost more than \$200,000 [9]. The cost of [7] can exceed \$10,000 due to their higher-end mmWave devices.
@@
As mentioned before, the proposed testbed can be operated autonomously, thanks to the “self-driving” capability enabled by the Donkey car [10]. It utilizes the front camera to identify the edges of the road and continuously control the steering angles and throttles. In the lab, we have setup a large canvas with dark tapes acting as the “road”. Fig. 2 shows the prototype of HawkRover.
@@ III-B Beam prediction with multimodal sensor fusion
The aim is to directly predict the best pair of codebooks from multisensor input, skipping channel estimation, as well as exhaustive search stage. Since enormous datasets can be generated from the HawkRover platform, our specific goal is to directly find the mapping:
$(b_{UE}^{*},b_{BS}^{*})=\max_{b_{UE},b_{BS}}f(\mathcal{C,L,I,P}),$ (1)
@@
In Fig. 5, the overall structure of the fusion algorithm is presented. At the high level, each sensor’s input is processed through convolutional neural network (CNN) and recurrent neural network (RNN). The former is usually used to process image, while the latter is for sequential data, perfectly fit for numerical wireless and location data. A fusion block is applied to concatenate each extracted feature, and then all features are fed into LSTM block for temporal correlations. Lastly, the softmax-based output layer is utilized to determine the beam index with highest probability.
In CNN, convolution (Conv), rectified linear unit (ReLU), and pooling are three fundamental building blocks. Each serves a distinct function. Conv layers perform a mathematical operation that takes two inputs, such as an image matrix and a filter or kernel, and produces a feature map. This operation involves sliding the filter over the image and computing the dot product at each position. Convolution preserves the spatial relationship between pixels by learning image features based on small squares of input data. The ReLU function applies a non-linear activation to the output of the convolution operation. It replaces all negative pixel values in the feature map with zero. This non-linearity introduces the ability to solve more complex problems by allowing the network to create complex mappings from the input to the output, improving the network’s ability to learn from the data. Pooling reduces the dimensionality of each feature map while retaining the most important information. Max pooling, for example, slides a window over the input and takes the maximum value at each step. By doing so, pooling layers reduce the number of parameters, lessen computational complexity, and provide a form of translation invariance. They help the network to be more robust to the exact location of features in the input space.
@@ IV A Case Study and Results
In this section, we provide a case study for one particular dataset collected. The location is a typical lab space. Note that during data collection, normal daily activities are not restricted, for example, people walking in the room, lab furniture movements, etc. This, in turn, shows the robustness and realism of the dataset.
@@
The collected data is divided into two parts: training (80%) and testing (20%). Both training process and testing experiments are implemented in PyTorch. As shown in Fig. 7, the optimal beamforming prediction result based on sensor fusion algorithm is presented. In DL, Top-$K$ is often used as an evaluation metric to assess the performance. For example, in our case, top-1 accuracy means the model’s most probable prediction must be correct for it to count as a success. Meanwhile, top-5 accuracy means any of the model’s five most probable predictions can be correct. We showcase the results from top-1 to top-5, each with three different settings, namely, LiDAR only, LiDAR + camera fusion, and LiDAR + Camera + IMU position fusion.
@@
We need to note that the mmWave spatial fingerprints of the lab room may have also contributed to this performance. As shown in our early works [12, 14, 13], mmWave signal can have a spatial resolution to sub-10 cm, and the DL may have learnt this implicit relationship of location and optimal beam directions. Nevertheless, sensor fusion approach can significantly reduce the beam tracking overhead, making it highly desirable feature for vehicular scenarios.
@@ V Conclusions
mmWave bands are emerging as a viable solution for CAV communication. However, their high attenuation and mobility management pose considerable challenges. Our study introduces HawkRover, an autonomous and cost-effective testbed for collecting comprehensive mmWave and sensor data such as from LiDAR, cameras, and IMUs, to enhance vehicular communication. By using DL-based sensor fusion, our method can predict signal paths more efficiently than iterative pilot signal processes. A case study for indoor scenario is given, which promises substantial progress in CAV communication research.